Consider the following data on U.S. GDP:



a. What was the growth rate of nominal GDP between 1994 and 2014? (Hint: The growth rate of a variable X over an N-year period is calculated as 100 ].)


b. What was the growth rate of the GDP deflator between 1994 and 2014?


c. What was real GDP in 1994 measured in 2009 prices?


d. What was real GDP in 2014 measured in 2009 prices?


e. What was the growth rate of real GDP between 1994 and 2014?


f. Was the growth rate of nominal GDP higher or lower than the growth rate of real GDP? Explain.

1
Expert's answer
2022-03-07T11:02:39-0500

a). The growth rate of nominal GDP for 20 years;

"=[{{(17,419-7,309)^\frac {1}{20}-1}}]\times 100"


"=(10,110^{0.05}-1)\times 100"


"=58.58\% \space or 59\%"


b). The growth rate of GDP deflator


"=[(108.3\\div 73.8) ^\\frac {1}{20}\n\n\u200b\n \n \u22121]\u00d7100"


"=(1.467^ \n{0.05}\n \u22121)\u00d7100"

"=0.76\%"


c). Real GDP in 1994 measured in 2009 prices?


"=7,309\div 73.8"


"= \$ \space99.04"


d). Real GDP in 2014 measured in 2009 prices?


"=17,419\div 108.3"


"=\$\space 160.84"


e). The growth rate of real GDP between 1994 and 2014?


"=[(160.84\\div 99.04) ^\\frac {1}{20}\n\n\u200b\n \n \u22121]\u00d7100"


"=2.45\%"


f).

The growth rate of the nominal GDP was higher than the growth rate of the real GDP since the Nominal GDP does not account for changes in price like inflation while The real GDP accounts for such changes like inflation hence, the growth rate tend to grow at a smaller speed due to rise in price.